How long will it take a runner to run 15km if the runner can maintain an average speed of 5m/s?


1
Expert's answer
2020-11-25T10:45:41-0500

We can find the time that the runner takes to run 15 km from the formula:


"t=\dfrac{s}{v}=\dfrac{15000\ m}{5\ \dfrac{m}{s}}=3000\ s=50\ min."

Answer:

"t=3000\ s=50\ min."